# Image Classification in High-Energy Physics: A Comprehensive Survey of Applications to Jet Analysis

### v2

@@ 7 Future direction and outlook
The future of ML and DL in HEP, particularly in jet analysis, is poised for transformative advancements. As researchers delve deeper into the petabyte-scale datasets generated by experiments like those at the LHC and QCD, the role of DL becomes increasingly vital. The potential implications of QML-baset Jet research for future particle physics experiments are significant. By demonstrating the effectiveness of QML for jet classification in section 5.1, this opens up new possibilities for improving the performance of particle physics experiments. Researchers could apply the suggested QML-based approaches to jet images to other HEP problems, such as signal versus background separation, anomaly detection, and particle track reconstruction. Furthermore, QML-based research on jet images could pave the way for the development of new quantum algorithms and hardware that could be used to solve complex problems in particle physics and other fields.
@@
There are multiple other compelling aspects and potential extensions that warrant further exploration, which are outlined here. For examples, for event-level analysis, a jet, in essence, cannot be entirely separated from an event’s remaining parts, yet "pure" jets can be achieved through grooming techniques. The utility of color connections is notable in various scenarios. The exploration into how to effectively demonstrate these effects is important, as there is potential in enhancing event-level analysis. The RNN approach, particularly RecNN, is easily adaptable for event-level analysis due to its natural fit into larger hierarchical structures. Previous studies have examined event analysis focusing solely on jets, utilizing simple RNN chains to reconstruct events from jets. When considering event-level implementation, structuring the entire event poses a significant challenge. Viewing each event as a structured data tree, where the entire event’s information is encapsulated in the nodes’ properties and their interconnections, is vital. Therefore, accurately representing each element and its connections within the event is crucial for developing neural network architectures. For Jet unsupervised learning, within the DNN framework, adjusting jet clustering could potentially enhance performance. Treating jet finding as a minimization problem presents an intriguing perspective, making it appealing to incorporate jet finding processes directly into event-level analysis. Another example, for new physics phenomena often display distinctive patterns related to their particle spectrum and decay modes. For instance, supersymmetry (SUSY) events typically generate a high number of final states, presenting a more complex hierarchical structure, and may include several soft leptons in electroweakino searches. Investigating whether DNNs can more effectively accommodate such topologies is also a worthwhile endeavor [64]. Moreover, distinguishing between quark-initiated and gluon-initiated Jets is crucial in collider experiments like the LHC. Discriminating between these Jets is challenging due to complex correlations in radiation patterns and non-perturbative effects like hadronization. AI methods, such as deep generative models, offer promising solutions to address this challenge [55]. Moving forward, there is a notable scarcity of published research on the application of auto-encoder (AE) for jet image processing, highlighting an opportunity for researchers to explore this field further. The potential for AE to significantly improve the separation of jet images from background noise presents a promising area of study. By focusing on this niche, researchers can contribute to advancing our understanding and methodologies in particle physics, potentially leading to more accurate and efficient analysis techniques.
@@
The complexity and volume of the data necessitate sophisticated analytical techniques that DL models, especially those based on CNNs and GNNs, are well-equipped to handle. These models excel in identifying intricate patterns and correlations within the data, making them invaluable for tasks such as jet tagging, particle tracking, and event classification. Furthermore, the scalability of DL models needs to be addressed to handle the increasing data rates from next-generation detectors and accelerators. Efficient training algorithms and model compression techniques will be essential for deploying these models in real-time analysis frameworks, enabling faster decision-making processes for data acquisition and retention. The future of DL in jet energy progression and estimation promises enhanced precision and efficiency. Innovations will likely focus on developing more sophisticated neural network models that can accurately predict jet energies in complex environments. Emphasis on real-time data analysis capabilities and integration with experimental workflows will be crucial, driving advancements in detecting and interpreting high-energy particle collisions more effectively and swiftly. The future of DL-based jet anomaly detection in HEP lies in advancing unsupervised learning techniques to uncover new physics signals hidden in complex data. Innovations in model interpretability and real-time processing will enhance detection capabilities. Cross-disciplinary collaboration will drive these advancements, leading to breakthroughs in identifying rare phenomena and expanding our understanding of the fundamental constituents of the universe. Others applications such as flavor tagging, pileup mitigation, and the reconstruction of decay chains. These DL-based Jet images can help in distinguishing between different types of particles based on their energy deposition patterns, aiding in the precise determination of particle origins and decay pathways. Additionally, they can be used for enhancing signal-to-noise ratios in complex collision environments, improving the accuracy of particle trajectory tracking, and in the analysis of jet substructure to identify specific decay processes, contributing to a deeper understanding of the underlying physics in high-energy collisions. The application of DL-based HEP jet images for tomography is promising. This approach has the potential to revolutionize how we visualize and analyze subatomic particles, offering unprecedented precision and insight. By leveraging DL techniques, researchers can improve the accuracy of tomographic reconstructions, enhancing our understanding of particle interactions and the fundamental structure of matter.
@@
Transfer learning (TL) is poised to revolutionize Jet HEP applications by leveraging pre-trained models from vast datasets to enhance performance on specific tasks with limited data. This approach can significantly reduce computational costs and training times, making it ideal for adapting models to new experiments or rare phenomena. As HEP experiments generate increasingly complex data, the ability to apply knowledge from one context to another will be invaluable for improving event classification, anomaly detection, and signal processing. Looking ahead, transfer learning will be crucial for efficiently extracting insights from new particle interactions and advancing our understanding of fundamental physics. Exploring advanced CNN architectures as source of prior knowledge like AlexNet, Fast-R-CNN, VGG, ResNet, GoogLeNet could offer potential improvements to target model that conduct many AI-based Jet tasks [41]. Generalizing the top tagger to classify other boosted objects, such as W/Z bosons, Higgs bosons, and BSM particles, is straightforward and extending it to partially-merged and fully resolved tops could enhance background rejection.
@@
Reinforcement learning (RL), with all its variants [81], in HEP jet applications is set to open novel pathways for optimizing experimental setups and data analysis strategies. By leveraging RL’s ability to learn optimal policies through interaction with an environment, future HEP experiments could see enhanced automation in event selection, detector alignment, and real-time data processing. The adaptability of RL models to dynamic systems makes them particularly suited for managing the complexities of particle collision events. As the technology matures, integrating RL into HEP could lead to significant advancements in experiment efficiency, discovery potential, and the ability to navigate vast datasets to uncover new physics phenomena. Additionally, federated learning (FL)-based computer vision [82] presents a promising frontier for Jet images applications, offering a pathway to harness collaborative model training while preserving data privacy and security. By distributing the learning process across multiple nodes, each holding its own subset of data, FL enables a collective improvement of models without direct data sharing. This approach is particularly suited for HEP collaborations spread across global institutions, where data locality and privacy concerns can limit traditional centralized training methods. Advancements in FL could lead to more robust, accurate models, enhancing our understanding of complex particle physics phenomena through cooperative, privacy-preserving analysis between different LHCs.
@@
The integration of large language models (LLMs) and generative AI into HEP has the potential to enhance the precision of particle detection and characterization. By leveraging these advanced AI models, researchers can identify subtle patterns and anomalies in jet images that might be missed by conventional methods. This improved accuracy is crucial for discovering new particles or interactions that could lead to breakthroughs in our understanding of the universe. For example, in the search for dark matter or other exotic particles, detecting faint signals amid a noisy background is a significant challenge. Generative AI can help by producing simulations that highlight these weak signals, allowing physicists to fine-tune their detection algorithms. Similarly, LLMs can assist by providing context and insight into these findings, suggesting potential theoretical implications and further areas of exploration. The application of LLMs and generative AI in HEP also promotes a more collaborative and interdisciplinary approach to research. By integrating AI experts with physicists, new methodologies and tools can be developed that leverage the strengths of both fields. This collaboration can lead to the creation of more sophisticated models that are specifically tailored to the needs of HEP. Furthermore, the insights gained from HEP research using AI can be applied to other fields, such as astrophysics, medical imaging, and materials science. This cross-pollination of ideas and techniques can drive innovation across multiple disciplines, leading to advancements that benefit a wide range of scientific endeavors.
@@ 8 Conclusion
Given the comprehensive assessment of ML and DL applications within the realm of HEP presented in this survey, it is evident that these techniques have significantly impacted various aspects of HEP experimentation and phenomenological studies. Through a detailed exploration of diverse DL approaches, including their application to HEP image classification, Jet particle analysis, and other pertinent areas, this paper has highlighted the potential of ML and DL techniques to enhance our understanding of particle physics phenomena. The analysis undertaken throughout this survey underscores the importance of leveraging AI models tailored to HEP images, as well as the significance of SOTA ML and DL techniques in advancing HEP inquiries. Specifically, the review has elucidated the implications of these techniques for tasks such as Jet tagging, Jet tracking, and particle classification, shedding light on their capabilities and limitations in addressing key challenges within the field. As we reflect on the current status of HEP grounded in DL methodologies, it becomes evident that while significant progress has been made, there remain inherent challenges that must be addressed to fully harness the potential of these approaches. These challenges include issues related to data quality, model interpretability, and generalization to diverse experimental conditions. Nonetheless, the survey also identifies promising avenues for future research endeavors, such as the development of novel DL architectures tailored to HEP data and the integration of domain-specific knowledge to enhance the performance of learning models. By addressing the challenges and leveraging the opportunities highlighted in this survey, researchers can continue to push the boundaries of HEP experimentation and pave the way for groundbreaking discoveries in particle physics using AI techniques.

### v3

@@ 5.4 Jet generation
5.5 Case studies in Jet tagging and classification
To provide a deeper insight into the applications of ML and DL techniques in jet classification for HEP, this section explores three critical case studies: top quark tagging, Higgs boson tagging, and photon jet classification.
  • Top quark tagging: This process is essential for distinguishing boosted top quarks from background events involving light quarks and gluons. Boosted top quarks often decay into a collimated spray of particles, which requires advanced tagging techniques to identify effectively. The ATLAS open data provides a comprehensive dataset for top quark tagging studies. Additionally, simulation tools like Delphes and MadGraph are frequently used to generate top quark events. Recent methods, including ParticleNet [49] and LorentzNet [31], have achieved significant improvements in classification accuracy by leveraging point-cloud representations of jets. These models employ graph-based architectures and permutation-invariant structures to enhance the discrimination power. Metrics such as classification accuracy and AUC have demonstrated significant improvements for top quark tagging using LorentzNet, achieving values exceeding 94% and 98.6%, respectively.
  • Higgs boson tagging: Is crucial for validating the SM and investigating potential new physics phenomena. Higgs bosons decaying into b-quarks generate jet structures with distinctive substructure features, making them a key focus for tagging studies. Datasets such as the CMS open data and the Higgs dataset from the University of California, Irvine ML repository serve as valuable resources for developing tagging algorithms. Traditional methods like BDTs and modern approaches such as CNNs have been employed extensively. Furthermore, advanced architectures like LGN [48] and ParticleNet [49] have demonstrated superior classification capabilities. By utilizing high-level kinematic features and DL techniques, classification accuracies exceeding 92% and AUC values surpassing 96% have been achieved, along with notable background suppression.
  • Photon Jet classification: Photon jet classification is a critical task for studying the quark-gluon plasma and distinguishing between direct photons and those originating from fragmentation processes. Quark-gluon datasets generated using PYTHIA8 simulations form the basis for training and evaluating classification models, with additional opportunities provided by CMS open data for analyzing real collision events. Advanced models such as EGNN [50] and PCT [51] have demonstrated effectiveness in capturing the energy deposits and angular distributions of particles within jets. Notably, state-of-the-art methods, such as EGNN, have demonstrated exceptional performance in photon jet classification tasks, achieving accuracies above 92% and AUC values exceeding 97%.
@@ 6 Future direction and outlook
The future of ML and DL in HEP, particularly in Jet analysis, is poised for transformative advancements. As researchers delve deeper into the petabyte-scale datasets generated by experiments like those at the LHC and QCD, the role of DL becomes increasingly vital. The potential implications of QML-baset Jet research for future particle physics experiments are significant. By demonstrating the effectiveness of QML for Jet classification in section 4.1, this opens up new possibilities for improving the performance of particle physics experiments. Researchers could apply the suggested QML-based approaches to Jet images and PCs to other HEP problems, such as signal versus background separation, anomaly detection, and particle track reconstruction. Furthermore, QML-based research on Jet tagging could pave the way for the development of new quantum algorithms and hardware that could be used to solve complex problems in particle physics and other fields.
@@
There are multiple other compelling aspects and potential extensions that warrant further exploration, which are outlined here. For examples, for event-level analysis, a Jet, in essence, cannot be entirely separated from an event’s remaining parts, yet "pure" Jets can be achieved through grooming techniques. The utility of color connections is notable in various scenarios. The exploration into how to effectively demonstrate these effects is important, as there is potential in enhancing event-level analysis. The RNN approach, particularly recursive neural network (RecNN), is easily adaptable for event-level analysis due to its natural fit into larger hierarchical structures. Previous studies have examined event analysis focusing solely on Jets, utilizing simple RNN chains to reconstruct events from Jets. When considering event-level implementation, structuring the entire event poses a significant challenge. Viewing each event as a structured data tree, where the entire event’s information is encapsulated in the nodes’ properties and their interconnections, is vital. Therefore, accurately representing each element and its connections within the event is crucial for developing neural network architectures. For Jet unsupervised learning, within the DNN framework, adjusting Jet clustering could potentially enhance performance. Treating Jet finding as a minimization problem presents an intriguing perspective, making it appealing to incorporate Jet finding processes directly into event-level analysis. Another example, for new physics phenomena often display distinctive patterns related to their particle spectrum and decay modes. For instance, supersymmetry (SUSY) events typically generate a high number of final states, presenting a more complex hierarchical structure, and may include several soft leptons in electroweakino searches. Investigating whether DNNs can more effectively accommodate such topologies is also a worthwhile endeavor [74]. Moreover, distinguishing between quark-initiated and gluon-initiated Jets is crucial in collider experiments like the LHC. Discriminating between these Jets is challenging due to complex correlations in radiation patterns and non-perturbative effects like hadronization. AI methods, such as deep generative models, offer promising solutions to address this challenge [63]. Moving forward, there is a notable scarcity of published research on the application of auto-encoder (AE) for Jet image processing, highlighting an opportunity for researchers to explore this field further. The potential for AE to significantly improve the separation of Jet images and PC from background noise presents a promising area of study. By focusing on this niche, researchers can contribute to advancing our understanding and methodologies in particle physics, potentially leading to more accurate and efficient analysis techniques.
@@
The complexity and volume of the data necessitate sophisticated analytical techniques that DL models, especially those based on CNNs and GNNs, are well-equipped to handle. These models excel in identifying intricate patterns and correlations within the data, making them invaluable for tasks such as Jet tagging, particle tracking, and event classification. Furthermore, the scalability of DL models needs to be addressed to handle the increasing data rates from next-generation detectors and accelerators. Efficient training algorithms and model compression techniques will be essential for deploying these models in real-time analysis frameworks, enabling faster decision-making processes for data acquisition and retention. The future of DL in Jet energy progression and estimation promises enhanced precision and efficiency. Innovations will likely focus on developing more sophisticated neural network models that can accurately predict Jet energies in complex environments. Emphasis on real-time data analysis capabilities and integration with experimental workflows will be crucial, driving advancements in detecting and interpreting high-energy particle collisions more effectively and swiftly. The future of DL-based Jet anomaly detection in HEP lies in advancing unsupervised learning techniques to uncover new physics signals hidden in complex data. Innovations in model interpretability and real-time processing will enhance detection capabilities. Cross-disciplinary collaboration will drive these advancements, leading to breakthroughs in identifying rare phenomena and expanding our understanding of the fundamental constituents of the universe. Others applications such as flavor tagging, pileup mitigation, and the reconstruction of decay chains. These DL-based Jet classification can help in distinguishing between different types of particles based on their energy deposition patterns, aiding in the precise determination of particle origins and decay pathways. Additionally, they can be used for enhancing signal-to-noise ratios in complex collision environments, improving the accuracy of particle trajectory tracking, and in the analysis of Jet substructure to identify specific decay processes, contributing to a deeper understanding of the underlying physics in high-energy collisions. The application of DL-based HEP Jet for tomography is promising. This approach has the potential to revolutionize how we visualize and analyze subatomic particles, offering unprecedented precision and insight. By leveraging DL techniques, researchers can improve the accuracy of tomographic reconstructions, enhancing our understanding of particle interactions and the fundamental structure of matter.
[…]
@@
TL, encompassing all its forms, including techniques like fine-tuning and domain adaptation [95, 96], is poised to revolutionize Jet HEP applications by leveraging pre-trained models from vast datasets to enhance performance on specific tasks with limited data. This approach can significantly reduce computational costs and training times, making it ideal for adapting models to new experiments or rare phenomena. As HEP experiments generate increasingly complex data, the ability to apply knowledge from one context to another will be invaluable for improving event classification, anomaly detection, and signal processing. Looking ahead, transfer learning (TL) will be crucial for efficiently extracting insights from new particle interactions and advancing our understanding of fundamental physics. Exploring advanced architectures as sources of prior knowledge, such as EfficientNet, vision Transformers (ViT), Swin Transformers, ConvNeXt, GNNs, neural ordinary differential equations (NODEs), physics-informed neural networks (PINNs), and and AutoML for architecture optimization, could offer substantial improvements to target models conducting AI-based Jet tasks [46]. These SOTA methods are better suited to handling the complexities of particle physics data compared to older architectures like AlexNet or VGG. Generalizing the top tagger to classify other boosted objects, such as W/Z bosons, Higgs bosons, and other particles, remains straightforward, and extending it to partially-merged and fully resolved tops could enhance background rejection.
Systematic errors are a significant concern in HEP experiments, particularly in image classification tasks involving jet analysis. These errors can arise from various sources, including detector calibration inaccuracies, biases in data reconstruction, and environmental factors during data acquisition. Addressing these uncertainties is crucial for the reliability and accuracy of ML models applied in HEP. One approach to mitigating systematic errors is through systematics-aware learning, which involves developing models that account for potential biases in the data. For instance, Estrade et al. [97] discussed the importance of creating benchmarks that capture realistic cases of systematic errors in HEP analysis to facilitate experimental comparisons of different techniques. Another strategy involves adversarial learning to eliminate systematic errors [98]. This paper discusses the application of adversarial domain adaptation in an unsupervised setting to reduce sample bias in supervised HEP event classifier training. The authors utilize a neural network with a gradient reversal layer to simultaneously enable signal versus background event classification while minimizing differences in the network’s response to background samples from different Monte Carlo models. Ghosh et al. [99] proposed classifiers that are fully aware of uncertainties and their corresponding nuisance parameters, demonstrating that this approach can enhance sensitivity to parameters of interest. By incorporating uncertainty directly into the learning process, models can achieve better performance compared to traditional strategies that do not account for such uncertainties.
To further enhance the mitigation of systematic errors, future research should focus on integrating uncertainty quantification and robust optimization directly into the design of ML architectures. This includes the development of hybrid models that combine traditional statistical techniques with modern ML approaches to explicitly model and correct for systematic effects. Additionally, employing advanced simulation techniques that better mimic real-world data will help reduce discrepancies between training datasets and experimental observations. Efforts should also be directed toward leveraging transfer learning to adapt models trained on simulated data to real-world experimental conditions more effectively. Another promising avenue is the application of federated learning in HEP, which enables collaborative training across multiple experimental datasets while preserving data privacy. This approach could be particularly effective in creating more generalized models that are less sensitive to dataset-specific biases. Finally, incorporating interpretability and explainability methods into systematic error analysis will help researchers better understand how models respond to uncertainties and biases, providing actionable insights to refine both experiments and ML methodologies. Such advancements will ultimately ensure that ML models in HEP are robust, transparent, and ready for real-world applications.
[…]
@@
RL, with all its variants [100, 101], in HEP Jet applications is set to open novel pathways for optimizing experimental setups and data analysis strategies. By leveraging reinforcement learning (RL)’s ability to learn optimal policies through interaction with an environment, future HEP experiments could see enhanced automation in event selection, detector alignment, and real-time data processing. The adaptability of RL models to dynamic systems makes them particularly suited for managing the complexities of particle collision events. As the technology matures, integrating RL into HEP could lead to significant advancements in experiment efficiency, discovery potential, and the ability to navigate vast datasets to uncover new physics phenomena. Additionally, federated learning (FL)-based computer vision [102] presents a promising frontier for Jet images applications, offering a pathway to harness collaborative model training while preserving data privacy and security. By distributing the learning process across multiple nodes, each holding its own subset of data, FL enables a collective improvement of models without direct data sharing. This approach is particularly suited for HEP collaborations spread across global institutions, where data locality and privacy concerns can limit traditional centralized training methods. Advancements in FL could lead to more robust, accurate models, enhancing our understanding of complex particle physics phenomena through cooperative, privacy-preserving analysis between different LHCs.
@@
The integration of large language models and generative AI [103] into HEP has the potential to enhance the precision of particle detection and characterization. By leveraging these advanced AI models, researchers can identify subtle patterns and anomalies in Jet tagging that might be missed by conventional methods. This improved accuracy is crucial for discovering new particles or interactions that could lead to breakthroughs in our understanding of the universe. For example, in the search for dark matter or other exotic particles, detecting faint signals amid a noisy background is a significant challenge. Generative AI can help by producing simulations that highlight these weak signals, allowing physicists to fine-tune their detection algorithms. Similarly, LLMs can assist by providing context and insight into these findings, suggesting potential theoretical implications and further areas of exploration. The application of LLMs and generative AI in HEP also promotes a more collaborative and interdisciplinary approach to research. By integrating AI experts with physicists, new methodologies and tools can be developed that leverage the strengths of both fields. This collaboration can lead to the creation of more sophisticated models that are specifically tailored to the needs of HEP. Furthermore, the insights gained from HEP research using AI can be applied to other fields, such as astrophysics, medical imaging, and materials science. This cross-pollination of ideas and techniques can drive innovation across multiple disciplines, leading to advancements that benefit a wide range of scientific endeavors.
@@ 7 Conclusion
Given the comprehensive assessment of ML and DL applications within the realm of HEP presented in this survey, it is evident that these techniques have significantly impacted various aspects of HEP experimentation and phenomenological studies. Through a detailed exploration of diverse DL approaches, including their application to HEP classification, Jet particle analysis, and other pertinent areas, this paper has highlighted the potential of ML and DL techniques to enhance our understanding of particle physics phenomena. The analysis undertaken throughout this survey underscores the importance of leveraging AI models tailored to HEP images and PCs, as well as the significance of SOTA ML and DL techniques in advancing HEP inquiries. Specifically, the review has elucidated the implications of these techniques for tasks such as Jet tagging, Jet tracking, and particle classification, shedding light on their capabilities and limitations in addressing key challenges within the field. As we reflect on the current status of HEP grounded in DL methodologies, it becomes evident that while significant progress has been made, there remain inherent challenges that must be addressed to fully harness the potential of these approaches. These challenges include issues related to data quality, model interpretability, and generalization to diverse experimental conditions. Nonetheless, the survey also identifies promising avenues for future research endeavors, such as the development of novel DL architectures tailored to HEP data and the integration of domain-specific knowledge to enhance the performance of learning models. By addressing the challenges and leveraging the opportunities highlighted in this survey, researchers can continue to push the boundaries of HEP experimentation and pave the way for groundbreaking discoveries in particle physics using AI techniques.
[…]
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
No data was used for the research described in the article.
Acknowledgement
The first author acknowledges that the study was partially funded by the Algerian Ministry of Higher Education and Scientific Research (Grant No. PRFU– A25N01UN2601 20230001).